Information for Contributors

The journal will consider for publication

Letters. Short summaries of original research, containing up to 2000 words, 4 figures, and 10 citations. Views. Invited mini-reviews of important research, usually no longer than 4000 words and containing no more than 8 figures and 25 citations. Reviews. Comprehensive overviews which summarize significant advances in a broad area of research, with up to 10000 words of text, 8 figures, and 100 citations.

Original papers and outlines for possible reviews should be submitted to the editor-in-chief. Send two copies with figures, one set of original figures, and one digital copy. Each paper will be assigned an action editor responsible for refereeing the paper. Authors may recommend one or more of the action editors. Accepted papers will appear with the name of the action editor who communicated the paper.

Digital submissions can be made by electronic mail or by IBM or Macintosh floppy disk. Whenever possible, include Postscript files for the figures.

The preferred format for papers is LaTEX, but manuscripts in troff are also acceptable. Papers in other formats will require more processing time.

There are no page charges for publication.

Please address submissions to:

Dr. Terrence Sejnowski The Salk Institute 10010 North Torrey Pines Road La Jolla, CA 92037

Electronic submissions:

terry@sdbio2.ucsd.edu